OpenStudy (anonymous):

Multiply (2+sqrt7)(4-sqrt7)

OpenStudy (anonymous):

(2+sqrt7)(4-sqrt7) =2(4)-2sqrt7+4sqrt7-(sqrt7)(sqrt7) =8+2sqrt7-7 =1+2sqrt7

OpenStudy (saifoo.khan):